OpenStudy (anonymous):

What is the solution to the equation -4x + 4 -2x = 2-(3x -2)?

OpenStudy (akashdeepdeb):

Open the brackets on the RHS -6x + 4 = 2 - 3x + 2 -3x = 4-4 = 0 x = 0 Understood? :)

OpenStudy (anonymous):

yes:-)